OpenStudy (anonymous):

medal If sine of x equals 1 over 2, what is cos(x) and tan(x)? Explain your steps in complete sentences.

OpenStudy (anonymous):

For what angle x does sin(x) = 1/2 ?

OpenStudy (anonymous):

It's a very commonly known value. Once you have the angle you will have all your answers.

OpenStudy (anonymous):

Hint: I would suggest you memorize what the sine and cosine of 30,45 and 60 degrees is because you'll see them a lot in exercises and tests.

OpenStudy (anonymous):

i think -30

OpenStudy (anonymous):

cos(x) = sqrt 3/2

OpenStudy (anonymous):

\[\sin x=\frac{ 1 }{ 2 }\] \[\sin ^2x+\cos ^2x=1,\cos ^2x=\frac{ 3 }{ 4 }\] \[\cos x=\pm \frac{ \sqrt{3} }{ 2 }\] \[\tan x=\pm ?\]

OpenStudy (anonymous):

is there a formula for tan

OpenStudy (anonymous):

\[tanx=\frac{ \sin x }{ \cos x }\]

OpenStudy (anonymous):

so 3/4/sqrt 3/2

OpenStudy (anonymous):

.61

OpenStudy (anonymous):

no \[\pm \frac{ \frac{ 1 }{ 2 } }{ \frac{ \sqrt{3} }{ 2 } }=\pm \frac{ 1 }{ \sqrt{3} }=\pm \frac{ \sqrt{3} }{ 3 } = \pm \frac{ 1.732 }{ 3 }=?\]

OpenStudy (anonymous):

thank

OpenStudy (anonymous):

Functions f(x) and g(x) are shown below: